OpenStudy (anonymous):

Write the quadratic function f(x) = x2 - 5x + 3 in vertex form.

Directrix (directrix):

Attached is the format for the Vertex form of a Quadratic. Source: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

Directrix (directrix):

Complete the Square: x^2 - 5x + 3 = x^2 - 5x + (-5/2)^2 + 3 - (-5/2)^2

Directrix (directrix):

x^2 - 5x + (-5/2)^2 + 3 - (-5/2)^2 [ ( x - 5/2) ( x - 5/2) ] + 3 - (25/4)

Directrix (directrix):

[ ( x - 5/2) ( x - 5/2) ] + 3 - (25/4) = ( x - 5/2 ) ^ 2 + 3 - 6.25 = ( x - 5/2 ) ^ 2 - 3.25

Directrix (directrix):

or, y = ( x - 2.25 ) ^ 2 - 3.25 @slevine In vertex form, the x-coordinate of the vertex is easy to see --> 2.25. The same holds for the y-coordinate of the vertex --> - 3.25

Directrix (directrix):

And, now something cool.

Directrix (directrix):

See the calculator at: http://tinyurl.com/6p8no9r